OpenStudy (noahbred):

Help? Solve. 4(a + 2) = 14 – 2(3 – 2a) A. no solution B. –2 C. all real numbers D. –1

OpenStudy (anonymous):

4(a + 2) = 14 - 2(3 - 2a) 4a + 8 = 14 - 6 + 4a by distribution 4a + 8 = 8 + 4a by combining common terms 4a = 4a by subtracting 8 a = a by dividing by 4 All Real Solutions.

OpenStudy (anonymous):

"all real numbers."

OpenStudy (anonymous):

thanks did i get it right?

OpenStudy (noahbred):

Um... idk didnt finifh test 2 questions left

OpenStudy (anonymous):

ahh okay need any more help?

OpenStudy (noahbred):

yes please

OpenStudy (noahbred):

Which equation would best help solve this problem? Five added to 2 times a number is equal to 7 less than 4 times the number. A. 5x + 2 = 7x – 4 B. 2x + 5 = 4x – 7 C. 4x +7 = 2x – 5 D. 2x – 5 = 4x + 7

OpenStudy (anonymous):

If the word number is replaced by x, and the words 'added', 'times', 'equal' and 'less than' are replaced by symbols, you would be left with: 5 + 2x = -7 + 4x If you slightly reorder is, you get: 2x + 5 = 4x - 7 The answer is b.

OpenStudy (anonymous):

is there anymore u need help with?

OpenStudy (anonymous):

u still there @noahbred

OpenStudy (noahbred):

Sorry i had to eat dinner sryyy

OpenStudy (noahbred):

Three added to 4 times a number is equal to 5 less than 6 times the number. What is the number? A. –8 B. –4 C. 4 D. 8

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

Three added to 4 times a number is equal to 5 less than 6 times the number. Three added to --> 3 + 4 times a number --> 4*x (x is the unknown number) is equal to ---> = 5 less than ---> -5 + (this one is tricky) 6 times the number. ---> 6x all together we have 3+ 4x = -5+6x

OpenStudy (noahbred):

so its 4

OpenStudy (anonymous):

yes that my answer

OpenStudy (noahbred):

100% thankss you play any video games?

OpenStudy (anonymous):

umm no not anymore